# Low bit-flip rate probabilistic error cancellation

### v1

@@ III.3 Computing Block-PEC coefficients
$=\sum_{\begin{subarray}{c}\mathsf{V}_{d}\in\mathbb{V}\end{subarray}}\alpha_{\mathsf{U}_{d}}(\mathsf{V}_{d})\sum_{\begin{subarray}{c}\mathsf{W}\in\mathbb{V}\\ \mathsf{V}^{\prime}_{d}\circ\mathsf{W}^{\prime}=\mathsf{V}^{\prime}\end{subarray}}\alpha_{\mathsf{U}_{d-1}\circ\cdots\circ\mathsf{U}_{1}}(\mathsf{W}^{\prime}),$ (24)
where $\mathsf{W}^{\prime}$ should be thought of as the result of commuting $\mathsf{V}_{1}$ to $\mathsf{V}_{d-1}$ through all gates and noise channels to the end of the circuit until reaching, but not crossing, the $d$-th layer.
@@
The effect of Block-PEC can be seen as replacing the relatively easy sampling method of the conventional PEC approach with a more complex sampling which extracts some classically computable quantities, and thereby reduces the number of quantum samples required for a given precision.
III.4 Hybrid Block-PEC for circuits with generic gates
We have previously established that $(\mathbb{U},\mathbb{V})$ is Pauli-$Z$ compatible, where $\mathbb{V}$ is the set of all Pauli-$Z$ string, and $\mathbb{U}=\mathbb{V}\cup\{\mathsf{Z}(\theta),\mathsf{ZZ}(\theta),\mathsf{CNOT}\}$. Both the T gate ($T=Z(\pi/4)$) and the phase gate ($S=Z(\pi/2)$) are in $\mathbb{U}$. The Hadamard gate $H$ together with the set $\mathbb{U}$ form a universal gate set, but $(\mathbb{U}\cup\{H\},\mathbb{V})$ is not Pauli-$Z$ compatible. This is especially problematic considering that the Hadamard gate cannot be replaced by a resource state in universal quantum computing jones2024 .
@@ III.4 Hybrid Block-PEC for circuits with generic gates
IV Performance analysis of Block-PEC
Section III derived a sampling cost improvement for block-PEC in the limiting case where additional controls are noiseless. In the following paragraphs, we examine more realistic situations for specific gate sets.
IV.1 Strictly bias-preserving gates
@@
Figure 2: Phase-flip propagation through the CNOT gate; the equality is between the unitaries associated to the circuits.
@@
Hence, for circuits with Toffoli gates, it will be necessary to resort to Hybrid Block-PEC described earlier. Nonetheless, the gates listed in Proposition IV.2 form a generating set for a rather large class of (sub)-circuits on which one can implement Block-PEC.
Figure 3: Performance of Block-PEC on randomly generated bias-preserving circuits, at $p=0.1$ uncorrelated dephasing noise; expressed as the sampling cost gain $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$ as a function of the qubit count $n$.
@@
Additionally, we studied analytically all combinations of two bias-preserving gates acting on at most 2 qubits, subjected to uncorrelated dephasing noise. We found out that the per-block method leads to an improvement when a CNOT follows a gate which acts non-trivially on the CNOT’s target qubit.
Proposition IV.3.
@@
The previous analysis can now be used to understand the possible gains for generic circuits that can be decomposed into sequences of sub-circuits with Pauli-$Z$ compatible ones interspersed with non Pauli-$Z$ compatible gates. Indeed, whenever the sub-circuits of the type $U_{(a)},U_{(b)},U_{(c)}$ appear within a circuit, Block-PEC yields an improvement over PEC (see Figure 5). The frequency of the appearance of such sub-circuits determines the efficiency of Block-PEC (see Figure 6 to see the improvement for randomly generated bias-preserving circuits). More concretely, if sub-circuits of type $U_{(a)},U_{(b)},U_{(c)}$ appear $k$ times throughout the circuit, the sampling cost decreases by a factor proportional to $kp^{2}$.
@@
We then numerically evaluated the sampling cost gain (defined as the ratio $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$). It shows that for a noise strength of $p=0.1$ and $n=8$, the sampling overhead of PEC is more than twelve times that of Block-PEC, as shown in Figure 3. As expected, whenever the strength of the noise decreases, the difference between the two methods vanishes. Indeed, for low noise strength, the probability of having more than one error per circuit becomes small enough so that the noise is well approximated by first sampling a single layer within the block and then sampling an error for this layer. This effectively means that sampling per layer or per block will be equivalent up to negligible factors and hence yield a similar sampling overhead.
Figure 5: Sampling cost gain $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$ for $U_{(x)}$ for $x\in\{a,b,c\}$, as a function of the uncorrelated phase-flip error $p$.
Figure 6: Performance of Block-PEC under uncorrelated noise for randomly generated circuits, made of Pauli-$Z$ compatible gates and Hadamard gates, in terms average sampling error rate (1000 samples per Monte Carlo simulation, with precision $\delta=0.03$), over 1000 runs.
IV.2 Partially bias-preserving circuits
@@
We use parity-preserving gates, and in particular RBS gates, as examples of partially bias-preserving gates. Other examples of such partially bias-preserving parity-preserving gates are the $\text{A}(\theta,\phi)$ gate (see e.g. Barkoutsos_2018 ; Gard_2020 ; vpe-vff ), the $\text{Givens}(\theta)$ gate (see e.g. Anselmetti_2021 ; Kivlichan_2018 ; arute2020 ), the $\text{XY}(\theta)$ gate (see e.g. Abrams_2020 ).
Theorem IV.5.
@@
Performance for RBS gate circuits.
@@
We have numerically evaluated the $8$-qubit pyramid represented in Figure 7 as it is a useful building block of several quantum machine learning algorithms (See Section V). Our simulations show that PEC requires more than 10 times as many quantum samples as Block-PEC when applied to a 8-qubit RBS diagonal (with randomly chosen RBS gate parameters) subjected to a phase-flip error probability of $p=0.1$ (see Figure 9), thereby offering sizable advantage in using our technique for high noise strength.
@@
V Block-PEC for concrete applications
In this section, we consider multiple applications in the context of computational finance (Section V.1) and machine learning (Section V.2), and which fit in the framework defined by Block-PEC.
All numerical simulations were implemented in Python, using Cirq cirq . We use the mitiq library mitiq to represent quasi-probabilistic distributions and execute the Monte Carlo simulations they correspond to. We implemented routines which pre-calculate the coefficients of Block-PEC’s quasi-probabilistic distributions, for circuits subjected to dephasing noise,
For convenience, dephasing noise will be modeled by uncorrelated dephasing noise channels, as described in Appendix B. In practical implementations on cat-qubit architectures, dephasing noise is correlated. In Appendix E, we show that for circuits subjected to correlated dephasing noise, Block-PEC yields the same performance gains witnessed in Section IV.1.
V.1 Mitigating errors in quantum option pricing
Option pricing is the task of estimating the payoff (value at expiration date) of option contracts. Building on quantum algorithms for amplitude estimation low-depth-qae , Stamatopoulos et al. defined a family of quantum Monte Carlo algorithms for option pricing Stamatopoulos_2020 . Such algorithms aim to calculate the expectation value of the payoff, by applying amplitude estimation to a unitary operator $\mathcal{A}$ such that
@@ V.1 Mitigating errors in quantum option pricing
We refer the interested reader to Stamatopoulos_2020 for a detailed explanation of the definition of the angles $\theta_{j}$ in order to compute the expectation value of an option with payoff $f$.
@@
Each implementation of $CY(\theta_{j})$ (as in Equation 42) contains one circuit pattern $U_{(a)}$ as $Y\left(\pm\frac{\theta_{j}}{2}\right)_{n}CNOT_{j-1,n}$. With $n-1$ repetitions of the circuit pattern $U_{(a)}$ in the implementation of $\mathcal{R}$, one can expect small gains when using Block-PEC as the sampling cost reduction for such sub-circuits is of lesser magnitude than for other types of sub-circuits.
@@
To quantify the efficiency of Block-PEC on this algorithm, we simulated the mitigation of errors on the payoff unitary $\mathcal{R}$. The results of our numerical simulations are aggregated in Figure 8. We have simulated the mitigation of errors on $4$-qubit payoff functions, implemented by the unitary $\mathcal{R}_{4}=Y(\theta_{0})CY(\theta_{1})_{0,4}\cdots CY(\theta_{4})_{3,4}$ for randomly chosen angles $\theta_{0},\cdots,\theta_{4}$. Taking $p=10^{-2}$, we computed the average error of PEC and Block-PEC over $1000$ samples. We observed that the sampling variance ratio is $\frac{\sigma^{2}_{\text{std}}}{\sigma^{2}_{\text{blk}}}\simeq 1.037548912$, which is in line with our expectation of a small sampling variance reduction when using Block-PEC over PEC.
@@
Figure 8: Average sampling error of PEC and Block-PEC on $4$-qubit payoff functions, at $p=0.01$ for uncorrelated noise; number of samples = 1000; sample variances: $\sigma^{2}_{\text{std}}\simeq 5.05e^{-6}$, $\sigma^{2}_{blk}\simeq 4.87e^{-6}$.
V.2 RBS-based parameterized quantum circuits
In the use of RBS gates, a common construction involves executing RBS gates for various angles following a pyramidal structure, as shown in Figure 7.
@@ V.2 RBS-based parameterized quantum circuits
For example, any orthogonal neural network can be mapped onto a pyramid of RBS gates acting on unary states onn . The training time of quantum (orthogonal) neural networks is in $O(1/\varepsilon^{2})$, where $\varepsilon$ is the accuracy of the estimation of the circuit’s output. Mitigating errors increases accuracy, and therefore reduces training time.
It is worth noting that using a unary encoding allows to detect errors by discarding non-unary output states, which can only appear when a bit flip occurred during the computation. While efficient, this is a potentially costly strategy. A low bit-flip error rate architecture makes the cost of this non-unary state error detection strategy negligible.
Figure 9: Performance of Block-PEC on diagonals and pyramids of RBS gates, at $p=0.1$ and $p=0.01$ respectively for uncorrelated noise; expressed as the sampling cost gain $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$ as a function of the qubit count $n$.
@@
The sampling cost improvement that Block-PEC yields for RBS-based parameterized quantum circuits depends on the particular architecture considered and the implementation of RBS gates. And since the calculation of the coefficients of the quasi-probabilistic distributions for Block-PEC does not depend on the angles of RBS gates, it is sufficient to calculate them once for every RBS-based parameterized circuit architecture, reducing the classical overhead of our Block-PEC approach.
This observation has multiple consequences. First, applying this reasoning to Noise-Extended PEC (NEPEC) nepec , which generalizes both PEC and Zero-Noise Extrapolation (ZNE), one can observe that mitigating errors by anti-diagonals is also an efficient strategy in the context of ZNE.
Secondly, anti-diagonals of RBS gates can be used to implement an unary data loader for the quantum option pricing algorithm, as presented in unary-option-pricing . Then Block-PEC yields an improvement over PEC on the entire execution of unary option pricing, and not just on the sub-circuit corresponding to the payoff function.
We observed that PEC requires more than twice as many quantum samples as Block-PEC when applied to a 8-qubit RBS pyramid subjected to a phase-flip error probability of $p=0.01$, as show in Figure 9.
@@
Since each implementation of $RBS(\theta)$ (as in Equation 42) contains one circuit pattern $U_{(b)}$ as $XCZ(\theta)_{j,k}CNOT_{j,k}$, one can expect noticeable gains when using Block-PEC. To quantify the efficiency of Block-PEC on this algorithm, we simulated the mitigation of errors on $4$-qubit RBS pyramids at $p=10^{-2}$. We computed the average error of PEC and Block-PEC over 1000 samples. We observed that the sampling variance ratio is $\frac{\sigma^{2}_{\text{std}}}{\sigma^{2}_{blk}}\simeq 1.684731146$, which suggests a noticeable improvement of the sampling variance when using Block-PEC (see Figure 10).
@@
Figure 10: Average sampling error of PEC and Block-PEC on $4$-qubit RBS pyramids at $p=10^{-2}$ for uncorrelated noise; number of samples = 1000; sample variances: $\sigma^{2}_{\text{std}}\simeq 3.51e^{-14}$, $\sigma^{2}_{blk}\simeq 2.08e^{-14}$.
VI Conclusion

### v2

@@ 3.3 Computing Block-PEC coefficients
$=\sum_{\begin{subarray}{c}\mathsf{V}_{d}\in\mathbb{V}\end{subarray}}\alpha_{\mathsf{U}_{d}}(\mathsf{V}_{d})\alpha_{\mathsf{U}_{d-1}\circ\cdots\circ\mathsf{U}_{1}}(\mathsf{W}^{\prime}),$ (24)
where $\mathsf{W}^{\prime}=\mathsf{V}_{d}^{-1}\circ\mathsf{V}^{\prime}$ should be thought of as the result of commuting $\mathsf{V}_{1}$ to $\mathsf{V}_{d-1}$ through all gates and noise channels to the end of the circuit until reaching, but not crossing, the $d$-th layer.
@@
The effect of Block-PEC can be seen as replacing the relatively easy sampling method of the conventional PEC approach with a more complex sampling which extracts some classically computable quantities and in turn reduces greatly the number of quantum samples required for a given precision.
3.4 Hybrid Block-PEC for circuits with generic gates
We have previously established that $(\mathbb{U},\mathbb{V})$ is Pauli-$Z$ compatible, where $\mathbb{V}$ is the set of all Pauli-$Z$ string, and $\mathbb{U}=\mathbb{V}\cup\{\mathsf{Z}(\theta),\mathsf{ZZ}(\theta),\mathsf{CNOT}\}$. Both the $T$ gate ($T=Z(\pi/4)$) and the phase gate ($S=Z(\pi/2)$) are in $\mathbb{U}$. The Hadamard gate $H$ together with the set $\mathbb{U}$ form a universal gate set, but $(\mathbb{U}\cup\{H\},\mathbb{V})$ is not Pauli-$Z$ compatible.
@@ 3.4 Hybrid Block-PEC for circuits with generic gates
4 Performance analysis of Block-PEC
Section 3 introduced the possibility of having up to an exponential gain in using Block-PEC as the number $d$ of consecutive layers of bias-preserving gates grows. Yet, to make such statement precise, it is necessary to define a gate-set for the bias-preserving gates, noise models and most importantly a class of circuits of interest for which to conduct the analysis of the sampling gain. Section 4.1 performs this analysis by first considering a single bias-preserving gate giving analytic performance guarantees, and then extending the analysis for several layers of bias-preserving circuits. Two cases are studied here, relatively shallow circuits with random arrangements of bias-preserving gates, and deeper circuits corresponding to $SWAP$ networks [24] that are used for superconducting NISQ architectures. Section 4.2 extends the latter to a wide class of circuits that can be directly used to construct applications. The impact on end-to-end applications is considered in Section 5.
4.1 Strictly bias-preserving gates
@@
Refer to caption Refer to caption
(a) (b)
Figure 2: Phase-flip propagation through the $CNOT$ gate; the equality is between the unitaries associated to the circuits. In (a) the phase-flip commutes as it affects the control of the $CNOT$ while in (b) it propagates from the target to the control.
@@
Hence, for circuits with Toffoli gates, it will be necessary to resort to Hybrid Block-PEC described earlier. Nonetheless, the gates listed in Proposition 4.2 form a generating set for a rather large class of (sub)-circuits on which one can implement Block-PEC.
@@
Additionally, we studied analytically all combinations of two bias-preserving gates acting on at most 2 qubits, subjected to uncorrelated dephasing noise.
We found out that the per-block method leads to an improvement when a $CNOT$ follows a gate which acts non-trivially on the $CNOT$’s target qubit. In such case, the block would be limited to the single $CNOT$ and the additional $Z$-controls of the preceeding gate.
Proposition 4.3.
@@
This analysis can now be used to understand the possible gains for generic circuits that can be decomposed into sequences of sub-circuits with Pauli-$Z$ compatible ones, interspersed with non Pauli-$Z$ compatible gates. Indeed, whenever the sub-circuits of the type $U_{(a)},U_{(b)},U_{(c)}$ appear within a larger circuit, Block-PEC yields an improvement over PEC (see Figure 4). The frequency of the appearance of such sub-circuits determines the efficiency of Block-PEC. More concretely, if sub-circuits of type $U_{(a)},U_{(b)},U_{(c)}$ appear $k$ times throughout the circuit, the sampling cost decreases by a factor proportional to $kp^{2}$.
Figure 4: Sampling cost gain $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$ for $U_{(x)}$ for $x\in\{a,b,c\}$, as a function of the phase-flip error $p$.
@@
We then numerically evaluated the sampling cost gain defined as the ratio $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$. It shows that for uncorrelated dephasing noise with strength $p=0.1$ and $n=8$, the sampling overhead of PEC is more than twelve times that of Block-PEC, as shown in Figure 5 (left). As expected, whenever the strength of the noise decreases, the difference between the two methods vanishes. This can be seen for $p=0.001$ in Figure 5 (right). Indeed, for low noise strength, the probability of having more than one error per circuit becomes small enough so that the noise is well approximated by a process that would affect a single location in the whole block. This effectively means that sampling per gate, per layer or per block will be equivalent up to negligible factors and hence yield a similar sampling overhead.
Refer to caption Refer to caption
Figure 5: Performance of Block-PEC on randomly generated bias-preserving circuits, for a phasing noise with error probability $p=0.1$ (left) and $p=0.001$ (right). The vertical axis shows the sampling cost gain $(\gamma_{\text{std}}/\gamma_{\text{blk}})^{2}$ as a function of the qubit count $n$. The depth of the circuit is $n+1$ making them relatively shallow and requiring higher error probabilities to exhibit a sizable gain when using Block-PEC.
Performance evaluation for $SWAP$ networks with bias-preserving gates.
An important class of circuits for superconducting devices in the NISQ regime is that of $SWAP$ networks [24]. These circuits allow to optimally swap qubit states so that any two qubit on a line become adjacent during one execution of the network. This then allows to perform a gate between them before putting them apart and allowing for other interactions. In effect, this solves the problem of long range interaction on architectures that have spatially fixed registers, though possibly requiring to repeat the $SWAP$ network to take into account non commuting interaction gates.
Now, when the $SWAP$ gates are compiled into 3 $CNOT$s, and if the gates performing the interactions are bias-preserving, the whole circuit is indeed bias-preserving as well. The depth and nature of the gates with many instances of pattern $U_{(b)}$ (Fig. 3-b) hint at the higer interest of implementing Block-PEC for these circuits.
This is demonstrated in Figure 6 where we have considered $SWAP$ networks of increasing width. Their depth was arbitrarily set to one times their width (left) and 3 times their width (right). The interactions before each swap are $ZZ(\theta)$ 2-qubit rotation gates with random $\theta$. There the exponential nature of the gain appears clearly and is confirmed by the total residual squarred error that it smaller for the exponential fit by a factor up to 6.12 when compared to the quadratic fit.
Refer to caption Refer to caption
Figure 6: Sampling gain for $SWAP$ networks and $ZZ(\theta)$ gates. The depth is equal to the width (left) or 3 times the width (right). The error model is independent uncorrelated single qubit dephasing noise for each wire touched by a gate. The dephasing probability is 0.001. The deeper the circuit, the more advantageous it is to perform Block-PEC, with already a factor 4 for 9 qubits only in the deeper circuit. The total squarred error for the exponential fit are 0.0035 (left) and 0.0756 (right) while it is 0.0083 (left) and 0.4623 (right) for the quadratic one. This entails the exponential fit a respectively 2.38 and 6.12 better precision than the quadratic one.
4.2 Partially bias-preserving circuits
@@
We use parity-preserving gates, and in particular $RBS$ gates, as examples of partially bias-preserving gates. Other examples of such partially bias-preserving parity-preserving gates are the $\text{A}(\theta,\phi)$ gate (see e.g. [4, 14, 13]), the $\text{Givens}(\theta)$ gate (see e.g. [2, 20, 3]), the $\text{XY}(\theta)$ gate (see e.g. [1]).
Theorem 4.5.
@@
Performance evaluation for $RBS$ gate circuits.
@@
We have numerically evaluated the pyramid circuit represented in Figure 7 as it is a useful building block of several quantum machine learning algorithms (See Section 5). Our simulations (Figure 8) show that Block-PEC exponentially reduces the number of samples necessary to achieve a given precision for the error mitigated value for circuits as the depth of the circuit increases. For instance, extrapolating our numerical simulations with an exponential fit let us expect a 12 times sampling gain for 30 qubits at $p=0.001$.
@@
Figure 8: Performance of Block-PEC on pyramids of $RBS$ gates (Figure 7), with a single-qubit phase-flip noise model with error probability $p=0.001$. The exponential fit exhibits a residual squared error of $8.07.10^{-6}$ while the quadratic fit is at $1.646.10^{-5}$. The best fit is $\approx 0.0018e^{0.2905n}+0.9962$ where $n$ is the number of qubits in the pyramid circuit. For $n=30$ the expected sampling gain is about 12 times.
Performance evaluation for $SWAP$ networks with $RBS$ gates.
As for the case of fully bias-preserving circuits, $SWAP$ networks [24] are useful to consider in the case of partially bias-preserving circuits. Here again, when the $SWAP$ gates are compiled into 3 $CNOT$s, with $RBS$ gates for the interaction gates, the whole circuit is partially bias-preserving.
The efficiency of Block-PEC is particularly stricking in Figure 9 where we have considered $SWAP$ networks of increasing width. Their depth was arbitrarily set to one times their width (left) and 3 times their width (right). The interactions between the qubits before each swaps are $RBS(\theta)$ gates with random $\theta$. Again the exponential nature of the gain appears clearly and is confirmed by the fact that the exponential fit exhibits a total squarred error smaller than that of the quadratic fit by a factor up to 16.04.
Refer to caption Refer to caption
Figure 9: Sampling gain for $SWAP$ networks and $RBS$ gates. The depth is equal to the width (left) or 3 times the width (right). The error model is independent uncorrelated single qubit dephasing noise for each wire touched by a gate. The dephasing probability is 0.001. The deeper the circuit, the more advantageous it is to perform Block-PEC, with already a factor 4 for 9 qubits only in the deeper circuit. The total squarred error for the exponential fit are 0.0004 (left) and 0.0340 (right) while it is 0.0030 (left) and 0.5452 (right) for the quadratic one. This entails the exponential fit a respectively 7.50 and 16.04 better precision than the quadratic one.
5 Block-PEC for concrete applications
In this section, we consider multiple applications in the context of computational finance (Section 5.1) and machine learning (Section 5.2), and which fit in the framework defined by Block-PEC.
All numerical simulations were implemented in Python, using Cirq [9]. We use the mitiq library [21] to represent quasi-probabilistic distributions and execute the Monte Carlo simulations they correspond to. We implemented routines which pre-calculate the coefficients of Block-PEC’s quasi-probabilistic distributions, for circuits subjected to dephasing noise,
For convenience, dephasing noise will be modeled by uncorrelated dephasing noise channels, as described in Appendix B. In practical implementations on cat-qubit architectures, dephasing noise is correlated. In Appendix E, we show that for circuits subjected to correlated dephasing noise, Block-PEC yields the same performance gains witnessed in Section 4.1.
5.1 Mitigating errors in quantum option pricing
Option pricing is the task of estimating the payoff (value at expiration date) of option contracts. Building on quantum algorithms for amplitude estimation [15], Stamatopoulos et al. defined a family of quantum Monte Carlo algorithms for option pricing [28]. Such algorithms aim to calculate the expectation value of the payoff, by applying amplitude estimation to a unitary operator $\mathcal{A}$ such that
@@ 5.1 Mitigating errors in quantum option pricing
We refer the interested reader to [28] for a detailed explanation of the definition of the angles $\theta_{j}$ in order to compute the expectation value of an option with payoff $f$.
@@
Each implementation of $CY(\theta_{j})$ given by Equation 42 contains one circuit pattern $U_{(a)}$ as $Y\left(\pm\frac{\theta_{j}}{2}\right)_{n}CNOT_{j-1,n}$. With $n-1$ repetitions of the circuit pattern $U_{(a)}$ in the implementation of $\mathcal{R}$, one can expect small gains when using Block-PEC as the sampling cost reduction for such sub-circuits is of lesser magnitude than for other types of sub-circuits.
@@
To quantify the efficiency of Block-PEC on this algorithm, we simulated the mitigation of errors on the payoff unitary $\mathcal{R}$. The results of our numerical simulations are aggregated in Figure 10. We have simulated the mitigation of errors on $4$-qubit payoff functions, where we chose to implement them using the unitary $\mathcal{R}_{4}=Y(\theta_{0})CY(\theta_{1})_{0,4}\cdots CY(\theta_{4})_{3,4}$ for randomly chosen angles $\theta_{0},\cdots,\theta_{4}$. Taking $p=10^{-2}$, we computed the average error of PEC and Block-PEC over $1000$ samples. We observed that the sampling variance ratio is $\frac{\sigma^{2}_{\text{std}}}{\sigma^{2}_{\text{blk}}}\simeq 1.0375$, which is in line with our expectation of a small sampling variance reduction when using Block-PEC over PEC.
@@
Figure 10: Average sampling error of PEC and Block-PEC on $4$-qubit payoff functions, at $p=0.01$; number of samples = 1000; sample variances: $\sigma^{2}_{\text{std}}\simeq 5.05e^{-6}$, $\sigma^{2}_{blk}\simeq 4.87e^{-6}$.
5.2 RBS-based parameterized quantum circuits
@@
In the use of $RBS$ gates, a common construction involves executing $RBS$ gates for various angles following a pyramidal structure, as shown in Figure 7. For example, any orthogonal neural network can be mapped onto a pyramid of $RBS$ gates acting on unary states [19]. The training time of quantum (orthogonal) neural networks is in $O(1/\varepsilon^{2})$, where $\varepsilon$ is the accuracy of the estimation of the circuit’s output. Mitigating errors increases accuracy, and therefore reduces training time.
@@
It is worth noting that using a unary encoding allows to detect errors by discarding non-unary output states, which can only appear when a bit flip occurred during the computation. While efficient, this is a potentially costly strategy. A low bit-flip error rate architecture makes the cost of this non-unary state error detection strategy negligible. The sampling cost improvement that Block-PEC yields for RBS-based parameterized quantum circuits depends on the particular architecture considered and the implementation of $RBS$ gates. And since the calculation of the coefficients of the quasi-probabilistic distributions for Block-PEC does not depend on the angles of $RBS$ gates, it is sufficient to calculate them once for every $RBS$-based parameterized circuit architecture, further reducing the classical overhead of our Block-PEC approach.
@@
We observed that PEC requires more than twice as many quantum samples as Block-PEC when applied to a 8-qubit $RBS$ pyramid subjected to a phase-flip error probability of $p=0.01$. Since each implementation of $RBS(\theta)$ given by Equation 42 contains one circuit pattern $U_{(b)}$ as $XCZ(\theta)_{j,k}CNOT_{j,k}$, one can expect noticeable gains when using Block-PEC. To quantify the efficiency of Block-PEC on this algorithm, we simulated the mitigation of errors on $4$-qubit $RBS$ pyramids at $p=10^{-2}$. We computed the average error of PEC and Block-PEC over 1000 samples. We observed that the sampling variance ratio is $\frac{\sigma^{2}_{\text{std}}}{\sigma^{2}_{blk}}\simeq 1.6847$, which suggests a noticeable improvement of the sampling variance when using Block-PEC (see Figure 11).
@@
Figure 11: Average sampling error of PEC and Block-PEC on $4$-qubit $RBS$ pyramids at $p=10^{-2}$; number of samples = 1000; sample variances: $\sigma^{2}_{\text{std}}\simeq 3.51e^{-14}$, $\sigma^{2}_{blk}\simeq 2.08e^{-14}$.
6 Conclusion